# Validity of relativistic hydrodynamics beyond local equilibrium

### v1

@@ V Generalisation to 3+13+1D
V.1 The linearized moment equations
Motivated by the moment evolution equations of the Boltzmann equation, we consider a generalized relaxation-type model,
$\left[\partial_{\tau}+\hat{\textbf{F}}\right]\vec{\rho}=-\hat{\textbf{C}}\vec{\rho}\,$, (72)
which preserves its essential features. We note that the structure of the linearized moment equations parallels that of the Boltzmann equation itself. This equation can be thought of as a moment version of the RTA equation with $\hat{{\mathbf{C}}}[\vec{\rho}]=\hat{{\mathbf{\Gamma}}}\vec{\rho}-\hat{{\mathbf{G}}}\vec{\rho}$. This has the added advantage of being able to incorporate multiple time scales and therefore more generalized collision kernels. This structure typically emerges in momentum dependent RTA models [49, 50, 45].
The spectrum of $\hat{\mathbf{C}}$ should be positive semi-definite much like that of the spectrum of the linearized Boltzmann collision kernel. Since $\hat{\mathbf{C}}$ acts on an infinite-dimensional vector space, its spectrum typically contains a continuous component. The kernel of $\hat{\mathbf{C}}$ comprises the five conserved moments associated with energy, momentum, and number conservation. Additionally, within the moment space, there is one more null vector corresponding to the equilibrium moments, satisfying $\hat{\mathbf{C}}[\vec{\rho}^{~\text{eq}}]=0$. When the system is truncated, the collision kernel $\hat{\mathbf{C}}_{N}$ becomes finite dimensional and its spectrum discrete. The operator $\hat{\mathbf{\Gamma}}$ is positive definite and invertible, as it corresponds to the collision frequency term projected into the moment space—generally non-zero. In the Anderson–Witting form of the RTA, $\hat{\mathbf{\Gamma}}$ reduces to $\hat{\mathbf{I}}/\tau_{R}$ with $\hat{{\mathbf{G}}}\vec{\rho}\to\hat{\mathbf{\Gamma}}\vec{\rho}^{~\text{eq}}$, with a single eigenmode and decay rate given by the relaxation time $\tau_{R}$.
Here $\mathbf{\hat{F}}$ is the moment space Liouville operator. The operator $\mathbf{\hat{F}}$ has a matrix structure [51, 52, 53],
@@ V.1 The linearized moment equations
where $\nabla^{\mu}$ are the commoving derivatives and $\mathbf{F}_{\mu}$, $\mathbf{A}$ are matrix(linear) operators that couples the moments.
V.2 Hydrodynamics in $3+1$ D
For simplicity, we limit our analysis to the RTA collision kernel, where $\hat{\boldsymbol{\nu}}{R}=\hat{\mathbf{I}}/\tau_{R}$. Most of the discussion in this section does not rely on this specific form of $\hat{\boldsymbol{\nu}}_{R}$; it only requires that the operator be invertible with positive definite eigenvalues. However, certain explicit results—such as the modifications to the transport coefficients—are derived under the assumption of the AW-RTA form.
@@ V.2 Hydrodynamics in $3+1$ D
V.2.1 Solutions to the moment equation
@@
We now consider the case where $\mathbf{F}{\mu}\nabla^{\mu},\hat{\boldsymbol{\nu}}_{R}]=0$ holds at all times. In this situation the dynamics simplify, closely resembling the familiar 0+10+1D case. In particular, we have $\hat{\boldsymbol{\nu}}_{c}=\hat{\boldsymbol{\nu}}_{R}$, and the propagator reduces to
@@
$\quad+\sum_{k=0}^{\infty}\frac{(-1)^{k},\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau).$ (89)
Thus, the analysis of the commuting case proceeds in direct analogy with the 0+10+1D result. The solution (V.2.1) in the commuting case is obtained by performing the coordinate transformation $\tau’,\tau]\to[\xi’,0]$ and expanding the integrand functions in a Taylor series in $\xi$’. This procedure relies on $\xi$’ being monotonic and commuting with $\hat{\mathcal{F}}$, allowing it to act as a scalar parameter in $\hat{\mathcal{F}}_{0,\xi’}$. In the non-commuting case, such a simplification is no longer possible, and the modification of transport coefficients cannot be expressed in the same compact form. However we can still get expressions for the gradient and non-perturbative decompositions.
@@
$-\int_{\tau_{0}}^{\tau}d\hat{\mathcal{W}}_{\tau,\tau^{\prime}}\hat{\boldsymbol{\nu}}^{-1}_{c}\partial_{\tau^{\prime}}\vec{\rho}_{c}^{~\text{eq}}(\tau^{\prime})\,.$ (90)
From this it follows that (Appendix (C.1)),
$\vec{\rho}(\tau)$ $=\sum_{k=0}^{\infty}\left[-\hat{\boldsymbol{\nu}}^{-1}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau)$
$+\hat{\mathcal{F}}_{\tau,\tau_{0}}\hat{\mathcal{W}}_{\tau,\tau_{0}}~\left(\vec{\rho}{(\tau_{0})}-\sum_{k=0}^{\infty}~\left[-\hat{\boldsymbol{\nu}}^{-1}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau_{0})\right)\,.$ (91)
@@
For a solution that is analytic in $\tau$, each gradient term, $\left[-\hat{\boldsymbol{\nu}}^{-1}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau_{0})$ in $\vec{\pi}^{G}(\tau_{0})$ can be written in terms of the corresponding terms $\left[-\hat{\boldsymbol{\nu}}^{-1}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau)$ via a Taylor expansion about $\tau$. The Taylor coefficients then gives the appropriate modifications to the transport coefficients to the gradient expansion. Deriving an explicit form analogous to the 0+10+1D solution (62) or the commuting case (V.2.1) is, however, algebraically involved. In Appendix (C.2), we present these expressions explicitly for the Anderson–Witting kernel under suitable simplifying assumptions.
VI Conclusion
@@
In this work, we showed that solutions of the Boltzmann equation and its simplified models share a universal structural form, decomposing into a Chapman–Enskog gradient contribution and an exponentially decaying non-perturbative modes. This separation reflects the presence of two fixed points of the collision dynamics, namely the collisionless regime and local thermal equilibrium, where the collision kernel vanishes. The gradient contribution describes the behavior near equilibrium, while the non-perturbative terms encodes the approach to the collisionless regime. A consistent description of collective dynamics in interacting systems requires interpolation between these two limits.
@@
Israel–Stewart theory and its generalizations were originally developed to resolve the causality issues inherent in the gradient expansion. These prescriptions promote non-equilibrium contributions to independent fields that relax to the equilibrium in finite time. We demonstrated that the resulting equations, constructed from the gradient expansion, are structurally similar to the exact evolution equations of the anisotropies, and thus naturally accommodate non-perturbative dynamics. Furthermore, we showed that free-streaming behavior can be effectively reproduced within the gradient expansion, provided that transport coefficients are appropriately modified at each order. In this way, our work confirms—and extends through purely analytical methods—the results reported in [18, 19, 20, 21, 22].
@@
Relativistic hydrodynamics therefore emerges as an interpolation between free streaming and collective dynamics, rather than a strict expansion around equilibrium. In the case of quark–gluon plasma simulations, transport coefficients are typically tuned to reproduce experimental observations. The ability of relativistic hydrodynamics to connect the early-time, weakly interacting regime with the late-time equilibrium regime may therefore underlie its remarkable success in describing far-from-equilibrium systems such as the quark–gluon plasma.
VII Acknowledgments
@@ Appendix A General Formulas
$\rho_{n,l}^{eq}$ $=\int\frac{d^{3}p}{(2\pi)^{3}}(p^{0})^{n}\left(\frac{p^{z}}{p^{0}}\right)^{2l}f^{eq}$ (94)
@@
$\rho_{n,l}^{eq}(T,z)=\frac{T^{n+3}}{2\pi^{2}}\frac{G_{n+3,l}(z)}{(2l+1)}e^{\mu/T}$ (101)

### v2

@@ V Generalisation to 3+13+1D
V.1 The linearised moment equations
Motivated by the moment evolution equations of the Boltzmann equation [38, 53] , we consider a generalized relaxation-type model,
$\left[\partial_{\tau}+\hat{\textbf{F}}\right]\vec{\rho}=-\hat{\textbf{C}}\left[\vec{\rho}-\vec{\rho}^{~\text{eq}}\right]\,$, (73)
with $\partial_{\tau}\equiv u^{\mu}\partial_{\mu}$. We note that the structure of the linearized moment equations parallels that of the Boltzmann equation itself. This equation can be thought of as a moment version of the RTA Boltzmann equation with $\hat{{\mathbf{C}}}[\vec{\rho}]\equiv\hat{\boldsymbol{\nu}}_{R}\left[\vec{\rho}-\vec{\rho}^{~\text{eq}}\right]$ being a linearization of the Boltzmann collision kernel, typically around equilibrium. This has the added advantage of being able to incorporate multiple time scales and therefore more generalized collision kernels. This structure typically emerges in momentum dependent RTA models [54, 55, 49].
The spectrum of $\hat{\mathbf{C}}$ should be positive semi-definite much like that of the spectrum of the linearized Boltzmann collision kernel. Since $\hat{\mathbf{C}}$ acts on an infinite-dimensional vector space, its spectrum typically contains a continuous component. The kernel of $\hat{\mathbf{C}}$ comprises the five conserved moments associated with energy, momentum, and number conservation. The operator $\hat{\boldsymbol{C}}$ is positive semi-definite, as it corresponds to the collision frequency term projected into the moment space—generally non-zero. In the Anderson–Witting form of the RTA, $\hat{\boldsymbol{\nu}}_{R}$ reduces to $\hat{\mathbf{I}}/\tau_{R}$ with a single eigenmode and decay rate given by the relaxation time $\tau_{R}$ and the conservation laws being imposed via matchig conditions.
Here $\mathbf{\hat{F}}$ is the moment space Liouville operator. The operator $\mathbf{\hat{F}}$ has a the structure [56, 57, 58],
@@ V.1 The linearised moment equations
where $\nabla^{\mu}$ are the commoving derivatives and $\mathbf{F}_{\mu}$, $\mathbf{A}$ are matrix(linear) operators that couples the moments. Both $\partial_{\mu}$ and $\mathbf{\hat{F}}$ are dependent on the choice of frame i.e the definition of $u^{\mu}$.
V.2 Solution in $3+1$ D
For simplicity, we limit our analysis to the RTA collision kernel, where $\hat{\boldsymbol{\nu}}_{R}=\hat{\mathbf{I}}/\tau_{R}$. Most of the discussion in this section does not rely on this specific form of $\hat{\boldsymbol{\nu}}_{R}$; it only requires that the operator be invertible with positive definite eigenvalues. However, certain explicit results—such as the modifications to the transport coefficients—are derived under the assumption of the AW-RTA form.
@@ V.2 Solution in $3+1$ D
Solutions to the moment equation
@@
We now consider the case where $\mathbf{\hat{F}},\hat{\boldsymbol{\nu}}_{R}]=0$ holds at all times. In this situation the dynamics simplify, closely resembling the familiar 0+10+1D case. In particular, we have $\hat{\boldsymbol{\nu}}_{c}=\hat{\boldsymbol{\nu}}_{R}$, and the propagator reduces to
@@
$\quad+\sum_{k=0}^{\infty}\frac{\gamma(k+1,\xi_{0})}{k!}\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau).$ (90)
Thus, the analysis of the commuting case proceeds in direct analogy with the 0+10+1D result. The solution (V.2) in the commuting case is obtained by performing the coordinate transformation $\tau’,\tau]\to[\xi’,0]$ and expanding the integrand functions in a Taylor series in $\xi$’. This procedure relies on $\xi$’ being monotonic and commuting with $\hat{\mathcal{F}}$, allowing it to act as a scalar parameter in $\hat{\mathcal{F}}_{0,\xi’}$. In the non-commuting case, such a simplification is no longer possible, and the modification of transport coefficients cannot be expressed in the same compact form. However we can still get expressions for the gradient and non-perturbative decompositions.
@@
$-\int_{\tau_{0}}^{\tau}d\hat{\mathcal{W}}_{\tau,\tau^{\prime}}\hat{\boldsymbol{\nu}}^{-1}_{c}\partial_{\tau^{\prime}}\vec{\rho}_{c}^{~\text{eq}}(\tau^{\prime})\,$ (91)
where we have assumed $\hat{\boldsymbol{\nu}}$ is invertible. This can always be done by projecting the moment equations to the subspace orthogonal to the kernel of $\hat{\boldsymbol{C}}$.
From this it follows that (Appendix (D.1)),
$\vec{\rho}(\tau)$ $=\sum_{k=0}^{\infty}\left[-\hat{\boldsymbol{\nu}}_{R}^{-1}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau)$
$+\hat{\mathcal{F}}_{\tau,\tau_{0}}\hat{\mathcal{W}}_{\tau,\tau_{0}}~\left(\vec{\rho}{(\tau_{0})}-\sum_{k=0}^{\infty}~\left[-\hat{\boldsymbol{\nu}}^{-1}_{R}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau_{0})\right)\,.$ (92)
@@
For a solution that is analytic in $\tau$, each gradient term, $\left[-\hat{\boldsymbol{\nu}}^{-1}_{R}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau_{0})$ in $\vec{\pi}^{G}(\tau_{0})$ can be written in terms of the corresponding terms $\left[-\hat{\boldsymbol{\nu}}^{-1}_{R}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau)$ via a Taylor expansion about $\tau$. The Taylor coefficients then gives the appropriate modifications to the gradient expansion. Deriving an explicit form analogous to the 0+10+1D solution (62) or the commuting case (V.2) is, however, algebraically involved. In Appendix (D.2), we present these expressions explicitly for the Anderson–Witting kernel under suitable simplifying assumptions.
V.3 Relativistic Hydrodynamics
In this section, we provide a heuristic derivation for hydrodynamic equations in the 3+13+1D setting from the moment equations and how the rescaled transport coefficients emerge. Consider the hierarchy of moment equations
$\partial_{\tau}\vec{\rho}=-\left[\hat{\boldsymbol{\nu}}_{R}+\mathbf{\hat{F}}\right]\vec{\rho}+\hat{\boldsymbol{\nu}}_{R}\vec{\rho}^{~eq}\,$, (94)
Let $\mathcal{M}$ be the vector space of all moments. We define the ’hydrodynamic’ subspace $\mathcal{H}_{\rm N}$ as a $\rm N$-dimensional subspace of $\mathcal{M}$ that includes, but is not restricted to, the moments, $T^{\mu\nu},~{\rm n^{\mu}}$. We assume that the operator $\hat{\boldsymbol{\nu}}_{R}$ depend only on the chosen hydrodynamic moments, while the equilibrium moment $\vec{\rho}^{~eq}$ depends on the equilibrium variables $u^{\mu},T,\mu$.
Let $\mathbf{\hat{P}}_{\mathcal{H}}$ be the projection operator that projects to $\mathcal{H}$. For $\vec{\rho}\in\mathcal{M}$, the hydrodynamic moments $\vec{\rho}_{\mathcal{H}}=\mathbf{\hat{P}}_{\mathcal{H}}\vec{\rho}~\in~\mathcal{H}_{\rm N}$ and non-hydrodynamic moments $\vec{\rho}_{\perp}=\left[\mathbf{\hat{I}}-\mathbf{\hat{P}}_{\mathcal{H}}\right]\vec{\rho}$$\in~\mathcal{H}^{\perp}$ form a complimentary subspaces in $\mathcal{M}=\mathcal{H}_{\rm N}\oplus\mathcal{H}^{\perp}$. We rewrite these equations for the anisotropies $\vec{\pi}$ and project the evolution equation Eq.(94) into the hydrodynamic subspace, we get the hydrodynamic equations,
$\partial_{\tau}\vec{\pi}_{\mathcal{H}}$ $=-\mathbf{\hat{P}}_{\mathcal{H}}\left[\hat{\boldsymbol{\nu}}_{R}+\mathbf{\hat{F}}\right]\vec{\pi}_{\mathcal{H}}$
$~~~~-\mathbf{\hat{P}}_{\mathcal{H}}\left[\hat{\boldsymbol{\nu}}_{R}+\mathbf{\hat{F}}\right]\vec{\pi}_{\perp}-\mathbf{\hat{P}}_{\mathcal{H}}~\hat{\boldsymbol{\rm D}}\vec{\rho}^{~eq}\,.$ (95)
Here, we assumed the projector is independent of spacetime. This can always be arranged with an appropriate choice of moment basis. To close the system of equations within the hydrodynamic subspace, we can use the solution obtained in Eq.(V.2) and project it to the non-hydrodynamic subspace,
$\vec{\pi}_{\perp}$ $=\vec{\pi}^{~G}_{\perp}+\vec{\pi}^{~T}_{\perp}\,$, (96)
where $\vec{\pi}^{~G}_{\perp}$ is the projection of the gradient term and $\vec{\pi}^{~T}_{\perp}$ is the projection of the transient term. The gradient term $\vec{\pi}^{~G}_{\perp}$ can be written in term of the hydrodynamic moments as they are made up of equilibrium gradients, $\left[-\hat{\boldsymbol{\nu}}^{-1}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau)$. Specifically, it is possible to write ,
$\vec{\pi}_{\perp}=\mathbf{\hat{\Delta}}\left(\vec{\rho}_{\mathcal{H}}\right)\vec{\pi}_{\mathcal{H}}\,.$ (97)
which gives the closure scheme. Note that this does not contradict vector space linear independence; rather, the dynamic is such that the coefficients in the non-hydrodynamic moment basis can be written as functions of the coefficients of the hydrodynamic moment basis (See appendix D.3).
The initial condition term $\vec{\pi}^{~T}_{\perp}$, depends both on the hydrodynamic terms via $\hat{\mathcal{F}}_{\tau,\tau_{0}}\hat{\mathcal{W}}_{\tau,\tau_{0}}$ ( Eq. (93) ) and on initial distribution. We can write this contribution as
$\mathbf{\hat{P}}_{\mathcal{H}}\left[\hat{\boldsymbol{\nu}}_{R}+\mathbf{\hat{F}}\right]\vec{\pi}^{~T}_{\perp}=\vec{\beta}^{~\rm T}\left(\vec{\rho}_{\mathcal{H}},\tau_{0}\right)\,.$ (98)
We can then write the closed set of hydrodynamic equations,
$\partial_{\tau}\vec{\pi}_{\mathcal{H}}$ $=-\mathbf{\hat{P}}_{\mathcal{H}}\left[\hat{\boldsymbol{\nu}}_{R}+\mathbf{\hat{F}}\right]\left(\mathbf{\hat{I}}+\mathbf{\hat{\Delta}}\right)\vec{\pi}_{\mathcal{H}}$
$~~~~-\mathbf{\hat{P}}_{\mathcal{H}}~\hat{\boldsymbol{\rm D}}\vec{\rho}^{~eq}-\vec{\beta}^{~\rm T}\left(\vec{\rho}_{\mathcal{H}},\tau_{0}\right)\,.$ (99)
The corrections to the transport coefficients comes both from $\mathbf{\hat{\Delta}}$ and $\vec{\beta}^{~\rm T}$ terms.
VI Discussion
Different hydrodynamic models correspond to different choices of the hydrodynamic subspace $\mathcal{H}_{N}$ and a corresponding closure scheme. For ideal hydrodynamics, the hydrodynamic subspace contains only moments corresponding to $\varepsilon,~n,~u^{\mu}$ with a five dimensional hydrodynamic subspace $\mathcal{H}_{5}$. Along with the equation of state, these form a closed set of equations where the assumption of local thermal equilibrium gives the closure scheme, $\vec{\pi}=\vec{0}$.
Navier-Stokes theory employs the same hydrodynamic subspace $\mathcal{H}_{5}$, but adopts a first-order gradient closure scheme in which higher-order moments are set to their first order asymptotic values (see Eq. (7)). Higher-order theories such as the Burnett and super-Burnett formulations introduce additional gradient corrections to this closure, while remaining asymptotic in nature.
Figure 2: Relativistic hydrodynamic evolution $\sim$ gradient expansion with modified transport coefficients
In second-order Israel-Stewart type theories enlarge the hydrodynamic subspace is enlarged, introducing new variables $\pi^{\mu\nu},~\Pi,~q^{\mu},~V^{\mu}$ with their own dynamical equations. After a choice of frame, this adds an extra nine independent moments resulting in a fourteen-dimensional hydrodynamic subspace $\mathcal{H}_{19}$. The closure scheme again sets the higher-order moments to their asymptotic values. However, the new scheme is fundamentally different from the Navier-Stokes expansions in that the Israel-Stewart formulation incorporates non-asymptotic corrections. The resulting equations are not purely an expansion around equilibrium; in addition to introducing finite relaxation times for the anisotropic moments, they also allow for the description of free streaming dynamics. This is schematically illustrated in Fig. 2. While Israel-Stewart theories were originally introduced to cure acausality and instability, their extended dynamical structure intrinsically goes beyond a purely asymptotic expansion around local equilibrium.
Extending the hydrodynamic subspace to $N>14$ with the asymptotic closure scheme for higher order moments allows a more accurate free streaming description. Near equilibrium, this will yield results similar to higher-order Chapman-Enskog type expansion. These two improvements address complementary regimes: enlarging the hydrodynamic subspace improves the description of far-from-equilibrium dynamics, while incorporating higher-order gradient corrections refines the near-equilibrium limit.
In section (V.3), we showed that the free streaming and relaxation type behaviour of higher order moments can be captured by altering the closure scheme to include transient contributions, rather than by introducing new dynamical variables. This provides a rigorous understanding of the renormalised transport coefficients proposed in previous studies [22, 25].
VII Conclusion
@@
In this work, we demonstrated that solutions of the Boltzmann equation and its simplified models share a universal structural form, decomposing into a Chapman-Enskog gradient contribution and an exponentially decaying non-perturbative mode. This separation reflects the presence of two fixed points of the collision dynamics, namely, the collisionless regime and local thermal equilibrium, where the collision kernel vanishes. The gradient contribution describes the behaviour near equilibrium, while the non-perturbative terms encode the approach to the collisionless regime. A consistent description of collective dynamics in interacting systems, therefore, requires interpolation between these two limits.
@@
Israel–Stewart theory and its generalisations were originally developed to resolve the causality issues inherent in the gradient expansion. These prescriptions promote non-equilibrium contributions to independent fields that relax to equilibrium on finite time scales. We showed that the resulting equations, constructed from the gradient expansion, are structurally similar to the exact evolution equations governing momentum anisotropies, and thus naturally accommodate non-perturbative dynamics. Furthermore, we demonstrated that free-streaming behaviour can be effectively reproduced within the gradient expansion, provided that transport coefficients are appropriately modified at each order. In this way, our work confirms and extends through purely analytical methods the results reported in [22, 23, 24, 25, 26].
@@
Relativistic hydrodynamics, therefore, emerges as an interpolation between free-streaming and collective dynamics, rather than as a strict expansion around equilibrium. In the context of quark–gluon plasma simulations, transport coefficients are typically tuned to reproduce experimental observations. The ability of relativistic hydrodynamics to connect the early-time, weakly interacting regime with the late-time equilibrium regime may therefore underlie its remarkable success in describing far-from-equilibrium systems such as the quark-gluon plasma.
VIII Acknowledgments
@@ Appendix A General Formulas
$\rho_{n,l}^{\rm eq}$ $=\int\frac{d^{3}p}{(2\pi)^{3}}(p^{0})^{n}\left(\frac{p^{z}}{p^{0}}\right)^{2l}f^{\rm eq}$ (101)
@@
$\rho_{n,l}^{\rm eq}(T,z)=\frac{T^{n+3}}{2\pi^{2}}\frac{G_{n+3,l}(z)}{(2l+1)}e^{\mu/T}$ (108)
We have
$\varepsilon$ $=\rho^{\rm eq}_{1,0}=\frac{T^{4}}{2\pi^{2}}G_{4,l}(z)e^{\mu/T}$ (109)
$P$ $=\rho^{\rm eq}_{1,1}=\frac{1}{3}\frac{T^{4}}{2\pi^{2}}G_{4,l}(z)e^{\mu/T}$ (110)
In general we have the relation,
$\rho^{\rm eq}_{1,l}$ $=\frac{1}{2l+1}\frac{T^{4}}{2\pi^{2}}G_{4,l}(z)e^{\mu/T}$ (111)
$=\frac{3}{(2l+1)}\frac{G_{4,l}(z)}{G_{4,1}(z)}P$ (112)
and the conformal limit,
$\rho^{\rm eq}_{1,l}=\frac{3}{(2l+1)}P$ (113)